The volume of a small piece of metal is 5.000cm3 at 200C and 5.014cm3 at 1000C. Determine the cubic expansivity of the metal?
1
Expert's answer
2020-06-01T14:24:16-0400

Solution.

"V_1=5.000cm^3;"

"T_1=200^oC;"

"V_2=5.014cm^3;"

"T_2=1000^oC;"

"\alpha_V-?;"

"\alpha_V=\dfrac{1}{V}\sdot\dfrac{\Delta V}{\Delta T};"


"\alpha_V=\dfrac{1}{5.000cm^2}\sdot\dfrac{0.014cm^2}{800^oC}=3.5\sdot10^{-6}1/^oC;"


Answer: "\alpha_V=3.5\sdot10^{-6}" oC-1.